OpenStudy (anonymous):

what is the equation of the parabola in vertex form? 0=y^2-x+2y-3

terenzreignz (terenzreignz):

Put the x on the left side first... \[\Large x = y^2 + 2y -3\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so (x+3)=(y+1)^2

terenzreignz (terenzreignz):

Oh... sure :Yeah, are you sure you needed help with this? LOL and now just bring the +3 back to the right side so that x is alone on one side.

OpenStudy (anonymous):

ok thanks

terenzreignz (terenzreignz):

Just a reminder that that +3 becomes -3 when you bring it over to the right side :)